There is quite a pile of playing cards sitting at my house.
A discovery I have made is as follows.
if 2 people are playing cards one card will be leftover.
if 3 people are playing cards 2 cards will be leftover.
if 4 people are playing cards 3 cards will be leftover .
if 5 people are playing cards for cards will be leftover.
if 6 people are playing cards 5 cards will be leftover.
if seven people are playing cards no cards will be leftover
if I have the minimum amount of cards to satisfy the above conditions how many cards do I have?

Answer by Glaviolette(140)

Since there are no cards left over when 5 or 7 people are playing, that means the number of cards is divisible by 5 and 7.
5 * 7 = 35
35/2 = 17 r 1
35/3 = 11 r 2
35/4 = 8 r 3
35/6 = 5 r 5
Therefore, 35 is the minimum number of cards possible.